Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Yi Qiang
Log entry time 08:37:11 on December 09, 2008
Entry number 252683
This entry is a followup to: 252552
keyword=Experiment Summary for Dec 8, Owl - Swing
Summary for Dec 8 (owl - swing)
Smooth data taking.
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/12/08 owl | 3.84 |
0.42 | 3.74 |
2008/12/08 day | 4.82 |
1.19 | 1.98 |
2008/12/08 swing | 2.75 |
2.08 | 3.18 |
total | 11.41 |
3.69 | 8.9 |
Main events/issues:
- 00:00-03:00 Ingector problem with a bad power supply.
- 14:50-15:10 EPR measurement
- 15:10-19:30 Moller measurement: 54%
- 21:50-23:00 Beam stops for fixing Pockell cell power supply problem.
- other time: production
Plans for the next few days:
- Production
- RF recovery
- Moller measurement
- Reference cell test